Week of 140915

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions to join the phone conference can be found here.

General Information

  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web

Monday

Attendance:

  • local: Maria Alandes (chair, minutes), Massimo Lamanna (CERN Storage), Maarten Litmaath (ALICE), Raja Nandakumar (LHCb), Tsung-Hsun Wu (ASGC)
  • remote: You-Jin Cha (KISTI), Michael Ernst (BNL), Jose Flix (PIC), Lisa Giacchetti (FNAL), Kyle Gross (OSG), Tiju Idiculla (RAL), Rolf Rumler (IN2P3), Onno Zweers (NL-T1)

Experiments round table:

  • ALICE -
    • NTR

  • LHCb reports (raw view) -
    • Mostly MC and User jobs over the week-end. Plan for re-stripping to start in October.
    • T0: NTR
    • T1: NTR
    • T2: Possible problem at NCBJ.pl, after downtimes - recurring network issues causing transfer problems out of NCBJ.
    • Services: NTR

Sites / Services round table:

  • ASGC: NTR
  • BNL: NTR
  • CNAF: Not present
  • FNAL: NTR
  • GridPP: Not present
  • IN2P3: A full day outage is scheduled for Tuesday next week to perform upgrade and reconfiguration for network setup. Affected services are Batch, dCache and DB. More details on the exact times on Wednesday next week.
  • JINR: Not present
  • KISTI: NTR
  • KIT: Not present
  • NDGF: Present but lost connection
  • NL-T1: NTR
  • OSG: NTR
  • PIC: NTR
  • RAL: NTR
  • RRC-KI: Not present
  • TRIUMF: Not present

  • CERN batch and grid services:
    • CERN CvmFS Tomorrow on Tuesday from 10:00 /cvmfs/lhcb.cern.ch and /cvmfs/cms.cern.ch stratum zeros will be migrated from 2.0 to 2.1. All mounting sites should be unaffected.
    • CERN FTS On Thursday this week the FTS fts3.cern.ch will be updated - should be transparent.
  • CERN storage services:
    • Castor CMS upgrade performed without any problems.
    • Next upgrades for the remaining experiments:
      • ATLAS: Tuesday from 9am to 10am
      • ALICE: Tuesday from 2pm to 3pm
      • LHCb: Thursday from 2pm to 3pm
    • Massimo clarifies to Raja that this upgrade is to version 2.1.14 which still includes xrootd 3. It's a minor upgrade that should be transparent.
  • Databases: Not present
  • GGUS: Not present
  • Grid Monitoring: Not present
  • MW Officer: Not present

AOB:

Thursday

Attendance:

  • local: Maria Alandes (chair, minutes), Nacho Barrientos (Grid&Batch), Marcin Blaszczyk (DB), Maria Dimou, Massimo Lamanna (CERN Storage), Maarten Litmaath (ALICE), Andrea Manzi (MW Officer), Tsung-Hsun Wu (ASGC)
  • remote: Dea-Ham Kim (KISTI), Jeremy Coles (GridPP), Michael Ernst (BNL), Jose Flix (PIC), Lisa Giacchetti (FNAL), Kyle Gross (OSG), John Kelly (RAL), Rolf Rumler (IN2P3), Dennis Van Dok (NL-T1), Sonia (CNAF), Pavel (KIT)

Experiments round table:

  • ALICE -
    • NTR

Sites / Services round table:

  • ASGC: There will be a downtime on 29th September to upgrade the MySQL DB used for the network link with Amsterdam.
  • BNL: NTR
  • CNAF: NTR
  • FNAL: NTR
  • GridPP: NTR
  • IN2P3: NTR
  • JINR: Not present
  • KISTI: NTR
  • KIT: NTR
  • NDGF: Not present in the Video Conference due to connection problems that will be reported to the Vydio team. They sent the report by mail:
    • Issues with Atlas jobs at UNICPH was tracked down, UCPH tier1 is back in production.
    • Downtimes at HPC2N,UiB, and NSC yesterday all finished succesfully.
    • There will be a 3-hour dCache pool maintenance at CSC on September 30.
    • Some Alice data unavailable.
  • NL-T1: There will be a major network upgrade on the week of 6th to 10th October. All systems are affected and may experience downtimes for short intervals of time. See GOCDB Downtime
  • OSG: NTR
  • PIC: NTR
  • RAL: NTR
  • RRC-KI: Not present
  • TRIUMF: Not present

  • CERN batch and grid services: NTR
  • CERN storage services: All Castor upgrades announced on Monday were performed and went OK.
  • Databases: Ongoing migration from Oracle Streams to Golden Gate replication technology:
    • Migrated LHCb Conditions DB without any problems
    • Tuesday next week: ATLAS to TRIUMF migration
    • Wednesday next week: ATLAS to RAL migration
  • GGUS: NTR
  • Grid Monitoring: Not present
  • MW Officer: NTR

AOB:

Topic attachments
I Attachment History Action Size Date Who Comment
Unknown file formatpptx 2014-09-16.pptx r1 manage 2861.3 K 2014-09-15 - 17:24 MariaDimou Final GGUS slides for the MB
Edit | Attach | Watch | Print version | History: r12 < r11 < r10 < r9 < r8 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r12 - 2014-09-18 - MariaALANDESPRADILLO
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback